OpenStudy (anonymous):

Which is the correct classification of this system of equations? -2x + y = 7 3y = 6x + 4 parallel coincident intersecting

OpenStudy (precal):

solve for y for both of them and then compare the slopes if slopes are the same it is parallel if lines are the same it is coincident example 2x+4y=8 is the same as x+2y=4 notice all they did was multiply a 2 to the second equation or I should say distributed a 2 because 2(x+2y=4) is the same as 2x+4y=8 if the slopes are different, then they probably intersect

OpenStudy (anonymous):

parallel

OpenStudy (anonymous):

parallel